# MaPPER: Multimodal Prior-guided Parameter Efficient Tuning for Referring Expression Comprehension

### v1

@@ Abstract
Referring Expression Comprehension (REC), which aims to ground a local visual region via natural language, is a task that heavily relies on multimodal alignment. Most existing methods utilize powerful pre-trained models to transfer visual/linguistic knowledge by full fine-tuning. However, full fine-tuning the entire backbone not only breaks the rich prior knowledge embedded in the pre-training, but also incurs significant computational costs. Motivated by the recent emergence of Parameter-Efficient Transfer Learning (PETL) methods, we aim to solve the REC task in an effective and efficient manner. Directly applying these PETL methods to the REC task is inappropriate, as they lack the specific-domain abilities for precise local visual perception and visual-language alignment. Therefore, we propose a novel framework of Multimodal Prior-guided Parameter Efficient Tuning, namely MaPPER. Specifically, MaPPER comprises Dynamic Prior Adapters guided by an aligned prior, and Local Convolution Adapters to extract precise local semantics for better visual perception. Moreover, the Prior-Guided Text module is proposed to further utilize the prior for facilitating the cross-modal alignment. Experimental results on three widely-used benchmarks demonstrate that MaPPER achieves the best accuracy compared to the full fine-tuning and other PETL methods with only 1.41% tunable backbone parameters.
@@
Ting Liu1,211footnotemark: 133footnotemark: 3, Zunnan Xu [^1] footnotemark: 1, Yue Hu2, Liangtao Shi3, Zhiqiang Wang4, Quanjun Yin [^2] footnotemark: 2 1Tsinghua University, 2National University of Defense Technology, 3Hefei University of Technology, 4iFLYTEK Research
**footnotetext: Equal contributionfootnotetext: Corresponding authorfootnotetext: Work done at Tsinghua University
@@ 1 Introduction
To address these issues, we shift our focus to Parameter-Efficient Transfer Learning (PETL) Lester et al. (2021); Chowdhury et al. (2023); Wang et al. (2023a). PETL methods like Adapter tuning and Prompt tuning provide efficient ways to utilize pre-trained models by adjusting a small set of parameters instead of fine-tuning the entire network Xin et al. (2024c). This approach saves computational resources while still competitive performance improvements. By integrating PETL techniques, we can enhance our models’ flexibility and efficiency in adapting to REC. However, we empirically find that directly using these PETL methods cannot achieve satisfactory results in REC (see Figure 1). We argue the main reasons are twofold: 1) the target objects that require attention in REC often occupy local regions of uncertain size in images, and most existing PETL methods lack the crucial ability to extract multi-scale local semantics for visual perception. 2) REC is a task that strongly relies on multimodal alignment, and language-oriented adapters are obviously deficient in aligning with visual information. Recently, PETL methods have also been introduced into REC tasks Xiao et al. (2024); Liu et al. (2024c). HiVG Xiao et al. (2024) adopts LoRA to fine-tune the frozen CLIP model, but it is not an efficient enough approach due to the heavy alignment design using cross-attention module. In contrast, DARA Liu et al. (2024c) is a lightweight method in PETL paradigm. However, DARA does not fully address the need for local visual adaptation in the referring expression comprehension, potentially compromising the model’s ability to capture fine-grained visual details.
@@ 2.1 Referring Expression Comprehension
Referring expression comprehension (REC) Yu et al. (2018); Yang et al. (2019); Deng et al. (2021); Xiao et al. (2023); Liu et al. (2024e); Xiao et al. (2024) aims to locate a local visual region in images by textual descriptions. Early propose-and-rank methods Liu et al. (2019b); Hong et al. (2019); Chen et al. (2019) follow a two-stage pipeline which first utilizes pre-trained object detectors to obtain a set of region proposals, which are then ranked based on their similarity scores with the given textual description. However, these two-stage methods face challenges in terms of the performance of the proposal generators and the additional ranking mechanisms. After the introduction of ViT, the Transformer-based methods Deng et al. (2021); Du et al. (2022); Yang et al. (2022); Zhu et al. (2022); Su et al. (2023); Liu et al. (2024c); Zhu et al. (2023) have recently emerged that significantly improve the grounding performance. Most recently, grounding multimodal large language models Li et al. (2023); Bai et al. (2023); Li et al. (2024); Wang et al. (2024, 2023b) have propelled the state-of-the-art (SOTA) performance, these works require a large amount of in-domain and other domain datasets. As REC models continue to scale up in size and complexity, fully fine-tuning becomes extremely high training cost.
@@ 2.2 Parameter-efficient Transfer Learning
The continuous expansion of pre-trained models demands significant computational resources and consumes considerable storage during fine-tuning Liu et al. (2024d). To address these challenges, researchers in the NLP and CV domain have explored PETL methods  Lester et al. (2021); Hu et al. (2022); Chen et al. (2022); Yuan et al. (2023); Zhou et al. (2024); Xu et al. (2024); Liu et al. (2024b). By focusing on updating only a small subset of parameters, PETL achieves a balance between maintaining high performance and ensuring computational efficiency. This method is particularly advantageous for deploying large-scale models, addressing the challenges posed by increasing model sizes while streamlining the adaptation process to new tasks. The main PETL methods can be classified into three categories: (i) selectively updating a tiny number of existing model parameters Guo et al. (2020); Zaken et al. (2021); (ii) adjusting newly added parameters to the model or its input Li and Liang (2021); Zhou et al. (2022); Xin et al. (2024b); (iii) applying low-rank factorization techniques to the parameters that require updates Hu et al. (2022); Karimi Mahabadi et al. (2021); Hao et al. (2023); Liu et al. (2024f); Xin et al. (2024a). Some pioneering works like ETRIS Xu et al. (2023) and DARA Liu et al. (2024c) sought to utilize adapters to adapt pre-trained models to referring image segmentation and referring expression comprehension, respectively. However, their proposed modules like Bridger Xu et al. (2023) and RA Liu et al. (2024c) are insufficient for capturing the complexity of multi-scale local visual features.

### v3

@@ Abstract
Referring Expression Comprehension (REC), which aims to ground a local visual region via natural language, is a task that heavily relies on multimodal alignment. Most existing methods utilize powerful pre-trained models to transfer visual/linguistic knowledge by full fine-tuning. However, full fine-tuning the entire backbone not only breaks the rich prior knowledge embedded in the pre-training, but also incurs significant computational costs. Motivated by the recent emergence of Parameter-Efficient Transfer Learning (PETL) methods, we aim to solve the REC task in an effective and efficient manner. Directly applying these PETL methods to the REC task is inappropriate, as they lack the specific-domain abilities for precise local visual perception and visual-language alignment. Therefore, we propose a novel framework of Multimodal Prior-guided Parameter Efficient Tuning, namely MaPPER. Specifically, MaPPER comprises Dynamic Prior Adapters guided by an aligned prior, and Local Convolution Adapters to extract precise local semantics for better visual perception. Moreover, the Prior-Guided Text module is proposed to further utilize the prior for facilitating the cross-modal alignment. Experimental results on three widely-used benchmarks demonstrate that MaPPER achieves the best accuracy compared to the full fine-tuning and other PETL methods with only 1.41% tunable backbone parameters. Our code is available at https://github.com/liuting20/MaPPER.
@@
Ting Liu [^1] footnotemark: 1, Zunnan Xu211footnotemark: 1, Yue Hu1, Liangtao Shi3, Zhiqiang Wang4, Quanjun Yin122footnotemark: 2 1College of Systems Engineering, National University of Defense Technology, 2Tsinghua University, 3Hefei University of Technology, 4iFLYTEK Research liuting20@nudt.edu.cn
**footnotetext: Equal contributionfootnotetext: Corresponding author
@@ 1 Introduction
To address these issues, we shift our focus to Parameter-Efficient Transfer Learning (PETL) Chowdhury et al. (2023); Wang et al. (2023a). PETL methods like Adapter tuning and Prompt tuning provide efficient ways to utilize pre-trained models by adjusting a small set of parameters instead of fine-tuning the entire network Xin et al. (2024c). This approach saves computational resources while still competitive performance improvements. By integrating PETL techniques, we can enhance our models’ flexibility and efficiency in adapting to REC. However, we empirically find that directly using these PETL methods cannot achieve satisfactory results in REC (see Figure 1). We argue the main reasons are twofold: 1) the target objects that require attention in REC often occupy local regions of uncertain size in images, and most existing PETL methods lack the crucial ability to extract multi-scale local semantics for visual perception. 2) REC is a task that strongly relies on multimodal alignment, and language-oriented adapters are obviously deficient in aligning with visual information. Recently, PETL methods have also been introduced into REC tasks Xiao et al. (2024); Liu et al. (2024c). HiVG Xiao et al. (2024) adopts LoRA to fine-tune the frozen CLIP model, but it is not an efficient enough approach due to the heavy alignment design using cross-attention module. In contrast, DARA Liu et al. (2024c) is a lightweight method in PETL paradigm. However, DARA does not fully address the need for local visual adaptation in the referring expression comprehension, potentially compromising the model’s ability to capture fine-grained visual details.
@@ 2.1 Referring Expression Comprehension
Referring expression comprehension (REC) Yu et al. (2018); Yang et al. (2019); Deng et al. (2021); Xiao et al. (2023); Liu et al. (2024e); Xiao et al. (2024) aims to locate a local visual region in images by textual descriptions. Early propose-and-rank methods Hong et al. (2019); Chen et al. (2019) follow a two-stage pipeline which first utilizes pre-trained object detectors to obtain a set of region proposals, which are then ranked based on their similarity scores with the given textual description. However, these two-stage methods face challenges in terms of the performance of the proposal generators and the additional ranking mechanisms. After the introduction of ViT, the Transformer-based methods Deng et al. (2021); Du et al. (2022); Yang et al. (2022); Zhu et al. (2022); Liu et al. (2024c); Zhu et al. (2023) have recently emerged that significantly improve the grounding performance. Most recently, grounding multimodal large language models Li et al. (2023); Wang et al. (2023b) have propelled the state-of-the-art (SOTA) performance, these works require a large amount of in-domain and other domain datasets. As REC models continue to scale up in size and complexity, fully fine-tuning becomes extremely high training cost.
@@ 2.2 Parameter-efficient Transfer Learning
The continuous expansion of pre-trained models demands significant computational resources and consumes considerable storage during fine-tuning Liu et al. (2024d). To address these challenges, researchers in the NLP and CV domain have explored PETL methods  Hu et al. (2022); Chen et al. (2022); Yuan et al. (2023); Liu et al. (2024b). By focusing on updating only a small subset of parameters, PETL achieves a balance between maintaining high performance and ensuring computational efficiency. This method is particularly advantageous for deploying large-scale models, addressing the challenges posed by increasing model sizes while streamlining the adaptation process to new tasks. The main PETL methods can be classified into three categories: (i) selectively updating a tiny number of existing model parameters Guo et al. (2020); Zaken et al. (2021); (ii) adjusting newly added parameters to the model or its input Li and Liang (2021); Zhou et al. (2022); Xin et al. (2024b); (iii) applying low-rank factorization techniques to the parameters that require updates Hu et al. (2022); Karimi Mahabadi et al. (2021); Hao et al. (2023); Liu et al. (2024f); Xin et al. (2024a). Some pioneering works like ETRIS Xu et al. (2023) and DARA Liu et al. (2024c) sought to utilize adapters to adapt pre-trained models to referring image segmentation and referring expression comprehension, respectively. However, their proposed modules like Bridger Xu et al. (2023) and RA Liu et al. (2024c) are insufficient for capturing the complexity of multi-scale local visual features.